# Beyond Convolution: A Taxonomy of Structured Operators for Learning-Based Image Processing

### 12 Mar 2026 v1

@@ 1 Introduction
The limitations of standard convolution are well illustrated by concrete tasks. In image denoising, applying the singular value decomposition (SVD) to local patches explicitly separates the structured (low-rank) signal component from the unstructured noise component [GZZ+14, CP22, HLL18]. In image super-resolution, interpolating bases such as B-splines encode smoothness priors that a fixed kernel cannot capture efficiently [DLH+16, SCH+16]. In classification, optimising the density function applied to the convolution kernel yields a spatially non-uniform weighting that improves convergence and accuracy [CP25c, CP25b]. In fuzzy signal processing, the F-transform with adaptive membership functions defines a projection operator whose bases are data-driven rather than fixed [CP24, PER06]. These examples suggest that the choice of operator is not merely an implementation detail but a fundamental modelling decision that encodes prior knowledge about the signal and the task.
@@
Several contributions in this taxonomy build upon and contextualise our own prior work. The decomposition-based family is grounded in our learning-based low-rank denoising approach [CP22], where a network predicts the optimal SVD thresholds for local image patches. The adaptive weighted family extends the framework of our weighted convolution with optimal density functions [CP25c, CP25b]. The basis-adaptive family encompasses our work on adaptive membership functions for the F-transform [CP24]. The computational aspects of operator optimisation are informed by our study of high-performance solvers for signal processing minimisation problems [CP25a]. Applications to medical imaging throughout the paper are motivated by our survey on ultrasound image processing [CP26].
@@ 4.1 Convolution with Density Functions
A density function $\bm{\Phi}\in\mathbb{R}^{K\times K}$ scales the kernel weights element-wise before applying the convolution. Given the standard kernel $\bm{w}$ and the density function $\bm{\Phi}$, the weighted convolution is defined as [CP25c]
@@
The optimal density function is computed by solving an outer optimisation problem [CP25c]: given a learning model $\mathcal{M}_{\bm{\Phi}}$ that minimises the loss function $\mathcal{L}$ with respect to the kernel weights $\bm{W}$ for a fixed $\bm{\Phi}$, the optimal density function $\bm{\Phi}^{*}$ is found by minimising the value of this minimised loss over all admissible density functions:
@@
Algorithm 2 Optimal density function computation [CP25c]
@@ 4.4 Properties and Task Suitability
Adaptive weighted operators are particularly effective in scenarios where the relative importance of neighbouring pixels varies across the image, such as tasks involving structured textures, edges, or anisotropic noise. Density function optimisation [CP25c, CP25b] improves convergence speed and final accuracy for both image-to-image (denoising, PSNR improvement of 6–7%) and image-to-label (classification, accuracy improvement of 7 percentage points) tasks, without increasing the number of trainable parameters. The computational overhead is modest ($\approx 7$% on GPU).
@@ 8.1 Structural Properties
Adapt. weighted (density) $\circ$ $\times$ $\circ$ $\mathcal{O}(K^{2})$ [CP25c]
@@ 8.3 Task Suitability Summary
For image-to-image tasks, decomposition-based and basis-adaptive operators are particularly well-suited because they explicitly encode structural properties of natural images (low-rank structure, multi-scale sparsity) that are relevant to denoising and super-resolution. For image-to-label tasks, adaptive weighted and attention-based operators are more appropriate, as they can capture global contextual information that is relevant for recognition and classification [CP25b]. In both cases, the optimal choice of operator depends on the specific signal statistics, the available training data, and the computational budget.
@@ 9 Open Challenges and Future Directions
All operators discussed in this paper are presented in 2D for clarity, but most have natural 3D extensions. For volumetric medical data (CT, MRI, ultrasound), structured operators that exploit the anisotropy of the acquisition process (e.g., different in-plane vs. through-plane resolution) could provide significant advantages [CP26]. Preliminary work on 3D weighted convolution is available at [CP25c].
@@
While the empirical performance of many operators has been demonstrated, their theoretical properties—approximation power, convergence of optimisation, generalisation bounds—are less well understood. Formal analysis of the optimisation landscape of the density function optimisation problem [CP25c] and the convergence of adaptive F-transform layers [CP24] are examples of the type of theoretical grounding that is needed across all operator families. Efficient solver design for the related minimisation problems also remains an open challenge [CP25a].
@@
The theoretical $50\%$ overhead of the weighted convolution is reduced to approximately $7\%$ on modern GPU hardware due to memory-level parallelism [CP25c]. More broadly, the computational properties of operators are increasingly architecture-dependent, motivating the design of operators that are co-optimised with the underlying hardware.

### 13 Mar 2026 v2

@@ 1 Introduction
The limitations of standard convolution are well illustrated by concrete tasks. In image denoising, applying the singular value decomposition (SVD) to local patches explicitly separates the structured (low-rank) signal component from the unstructured noise component [GZZ+14, CP22, HLL18]. In image super-resolution, interpolating bases such as B-splines encode smoothness priors that a fixed kernel cannot capture efficiently [DLH+16, SCH+16]. In classification, optimising the density function applied to the convolution kernel yields a spatially non-uniform weighting that improves convergence and accuracy [CP25b, CP25c]. In fuzzy signal processing, the F-transform with adaptive membership functions defines a projection operator whose bases are data-driven rather than fixed [CP24, PER06]. These examples suggest that the choice of operator is not merely an implementation detail but a fundamental modelling decision that encodes prior knowledge about the signal and the task.
@@
Several contributions in this taxonomy build upon and contextualise our own prior work. The decomposition-based family is grounded in our learning-based low-rank denoising approach [CP22], where a network predicts the optimal SVD thresholds for local image patches. The adaptive weighted family extends the framework of our weighted convolution with optimal density functions [CP25b, CP25c]. The basis-adaptive family encompasses our work on adaptive membership functions for the F-transform [CP24]. The computational aspects of operator optimisation are informed by our study of high-performance solvers for signal processing minimisation problems [CP25a]. Applications to medical imaging throughout the paper are motivated by our survey on ultrasound image processing [CP26].
@@ 4.1 Convolution with Density Functions
A density function $\bm{\Phi}\in\mathbb{R}^{K\times K}$ scales the kernel weights element-wise before applying the convolution. Given the standard kernel $\bm{w}$ and the density function $\bm{\Phi}$, the weighted convolution is defined as [CP25b]
@@
The optimal density function is computed by solving an outer optimisation problem [CP25b]: given a learning model $\mathcal{M}_{\bm{\Phi}}$ that minimises the loss function $\mathcal{L}$ with respect to the kernel weights $\bm{W}$ for a fixed $\bm{\Phi}$, the optimal density function $\bm{\Phi}^{*}$ is found by minimising the value of this minimised loss over all admissible density functions:
@@
Algorithm 2 Optimal density function computation [CP25b]
@@ 4.4 Properties and Task Suitability
Adaptive weighted operators are particularly effective in scenarios where the relative importance of neighbouring pixels varies across the image, such as tasks involving structured textures, edges, or anisotropic noise. Density function optimisation [CP25b, CP25c] improves convergence speed and final accuracy for both image-to-image (denoising, PSNR improvement of 6–7%) and image-to-label (classification, accuracy improvement of 7 percentage points) tasks, without increasing the number of trainable parameters. The computational overhead is modest ($\approx 7$% on GPU).
@@ 8.1 Structural Properties
Adapt. weighted (density) $\circ$ $\times$ $\circ$ $\mathcal{O}(K^{2})$ [CP25b]
@@ 8.3 Task Suitability Summary
For image-to-image tasks, decomposition-based and basis-adaptive operators are particularly well-suited because they explicitly encode structural properties of natural images (low-rank structure, multi-scale sparsity) that are relevant to denoising and super-resolution. For image-to-label tasks, adaptive weighted and attention-based operators are more appropriate, as they can capture global contextual information that is relevant for recognition and classification [CP25c]. In both cases, the optimal choice of operator depends on the specific signal statistics, the available training data, and the computational budget.
@@ 9 Open Challenges and Future Directions
All operators discussed in this paper are presented in 2D for clarity, but most have natural 3D extensions. For volumetric medical data (CT, MRI, ultrasound), structured operators that exploit the anisotropy of the acquisition process (e.g., different in-plane vs. through-plane resolution) could provide significant advantages [CP26]. Preliminary work on 3D weighted convolution is available at [CP25b].
@@
While the empirical performance of many operators has been demonstrated, their theoretical properties—approximation power, convergence of optimisation, generalisation bounds—are less well understood. Formal analysis of the optimisation landscape of the density function optimisation problem [CP25b] and the convergence of adaptive F-transform layers [CP24] are examples of the type of theoretical grounding that is needed across all operator families. Efficient solver design for the related minimisation problems also remains an open challenge [CP25a].
@@
The theoretical $50\%$ overhead of the weighted convolution is reduced to approximately $7\%$ on modern GPU hardware due to memory-level parallelism [CP25b]. More broadly, the computational properties of operators are increasingly architecture-dependent, motivating the design of operators that are co-optimised with the underlying hardware.